OpenStudy (anonymous):

simplify Y 1 -- (divied) -- 4 4 a y b 4y c 1 d 4

OpenStudy (anonymous):

\[\huge \frac{ \frac y4}{\frac 14}\] is this your equation...?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[\huge \frac y4 \div \frac 14\] \[=\huge \frac y4 \times \frac 41\]

OpenStudy (anonymous):

the bottom is choice though O_o

OpenStudy (anonymous):

isnt*

OpenStudy (anonymous):

sorry, now times it out. multiplying fractions, so times across (numerator x numerator over denominator x denominator)

OpenStudy (anonymous):

is it y 4y 1 4

OpenStudy (anonymous):

im not giving you the answer man, but i will help you work it out so it becomes: \[\huge \frac {y \times 4}{4 \times 1}\] can you solve from here?